### v1

# Exposing a Fatal Flaw in Sample-based Quantum Diagonalization Methods

Abstract
Quantum Selected Configuration Interaction (QSCI) methods (also known as Sample-based Quantum Diagonalization, SQD) have emerged as promising near-term approaches to solving the electronic Schrödinger equation with quantum computers. In this work, we show that QSCI methods face fundamental limitations that severely hinder their practical applicability. Using the nitrogen molecule and the iron-sulfur cluster [2Fe-2S] as examples, we demonstrate that while QSCI can, in principle, yield high-quality CI expansions similar to classical SCI heuristics in some cases, the method struggles with inefficiencies in finding new determinants as sampling repeatedly selects already seen configurations. This inefficiency becomes especially pronounced when targeting high-accuracy results or sampling from an approximate ansatz. In cases where the sampling problem is not present, the resulting CI expansions are less compact than those generated from classical heuristics, rendering QSCI an overall more expensive method. Our findings suggest a fatal flaw in QSCI methods as the inescapable trade-off between finding sufficiently many determinants and generating compact, accurate CI expansions. This ultimately hinders utility in quantum chemistry applications as QSCI falls behind more efficient classical counterparts.
[…]
1 Introduction
The full configuration interaction (FCI) method provides the exact solution to the electronic Schrödinger equation (within a given basis) and is thus considered the ground truth within quantum chemistry1, 2, 3. However, due to the exponential scaling in computing and memory requirements, exact FCI is limited to very small molecules. One of the largest demonstrations to date was by Gao et al. on propane (C3H8)/STO-3G4, who used a distributed FCI implementation on 256 nodes to treat a CI expansion of 1.3 trillion determinants. As a technical achievement, such work is certainly impressive. If one is willing to settle for near-exact FCI results (within, e.g., micro-Hartree precision), approximate methods such as selected configuration interaction (SCI) can obtain results much more affordably5. Various developments within electronic structure theory have focused on achieving near-FCI accuracy at a lower cost3, including methods based on the density matrix renormalization group (DMRG)6, many-body expanded FCI (MBE-FCI)7, 8, 9, and various flavors of SCI10, 11, 12, 13. Recently, such methods were put to the test in a “blind challenge” of predicting the ground-state energy of the benzene molecule (C6H6) in a cc-pVDZ basis set2. It revealed that a firm consensus between various near-exact FCI methods has still not been reached, even when agreement from calculations on smaller systems would have suggested so.
Applying quantum computers within the field of quantum chemistry has received significant attention in recent years14, 15, 16, 17. This is partly due to the expectation that quantum computers should be a natural candidate for solving the electronic Schrödinger equation as they inherently operate within an exponentially large Hilbert space. In the long term, provided fault-tolerant devices become a reality, quantum computers could, for example, be used to obtain ground state energies using quantum phase estimation18, 19. However, current near-term quantum computers are noisy and rather limited in terms of qubit count, error rates, and execution speed, which limits their practical advantage over classical computing architectures20. This has motivated the development of hybrid algorithms that take advantage of both quantum and classical computing architectures, resulting in shallower quantum circuits, which are more realistic options for practical quantum computing applications in the near term21, 22, 23. The variational quantum eigensolver (VQE) algorithm24, 25, 26 has emerged as a particularly popular near-term method. VQE-type methods rely on a parameterized ansatz, which is iteratively optimized using a hybrid quantum-classical feedback loop. The quantum device prepares a state and measures the expectation values over Pauli operators. The classical computer collects these measurements to evaluate expectation values over the fermionic Hamiltonian and updates the parameters of the parameterized ansatz using classical optimization methods. Although popular, VQE-type methods have several serious drawbacks related to the measurement process27 (which inherently has statistical noise) and the expressibility of the ansatz28, 29, 26.
The quantum selected configuration interaction (QSCI) method, introduced by Kanno et al.30, 31, addresses some of the challenges faced with VQE by drastically reducing the quantum workload, leveraging quantum devices solely for determinant selection. The central idea of the method is that some approximate ground state is prepared on a quantum computer, which is then used to draw samples corresponding to determinants in the computational basis. The CI Hamiltonian is then classically diagonalized in the basis of the selected determinants using conventional SCI diagonalization routines. We note that the classic diagonalization eventually becomes the computational bottleneck of the algorithm, provided enough determinants are discovered. Thus, the “quantum” part of QSCI lies solely in the selection of determinants for the SCI expansion. Such a division of work between quantum and classical resources has some clear advantages. First, since the quantum device is only used for selection, while energies are obtained classically, one removes the statistical errors associated with Hamiltonian measurement. Second, the CI vector is accessible on the classical device after the diagonalization, which allows computing expectation values over other quantum mechanical operators or correcting the energy of the variational wave function with Epstein-Nesbet perturbation theory32. In the original QSCI work, the dimension of the subspace (i.e., the number of determinants included in the CI expansion) was rather limited, with at most around a hundred determinants, in part due to a limited variational space when using rather small basis sets.
The QSCI method has also been demonstrated on a larger scale by Robledo-Moreno et al.33, under the name sample-based quantum diagonalization (SQD) with applications to potential energy curves of the nitrogen molecule (in a cc-pVDZ basis) and the iron-sulfur cluster model systems [2Fe-2S] and [4Fe-4S], with active spaces of (30e,20o) and (54e,36o), respectively. The demonstrations in Ref. 33 are relatively “large-scale” in nature, using CI expansion with millions of determinants. The central procedure is essentially identical to the original SCI paper but demonstrates the use of a large number of qubits (up to 77) and handling aspects such as treating quantum samples with incorrect particle numbers. The state preparation is handled with a local unitary cluster Jastrow ansatz (LUCJ)34, 35, with parameters assigned classically based on CCSD $t_{2}$-amplitudes. The SQD approach has since become a showcase for quantum-centric supercomputing36 and has been applied in several follow-up works, including the treatment of interaction energies with the potential energy surfaces of water and methane dimers37, spin state energetics of Fe(III)-NTA 38 and methylene39, the treatment of excited states40, and inclusion of environment effects through density matrix embedding theory41.
Separately, we note that an approach for determining the ground state of the SCI using quantum computers has also been proposed42. Unlike QSCI and SQD, determinants are selected using conventional SCI heuristics, and the focus is on leveraging variational quantum circuits to find the ground-state energy of the CI Hamiltonian.
The QSCI method introduces a potentially useful approach to determinant selection via quantum sampling, but several fundamental challenges limit its practical utility in real-world applications. First of all, the method arguably requires the preparation of a good state on the quantum device. Ideally, such a state should provide a probability distribution that closely matches the exact FCI solution since one would otherwise miss important determinants or include a significant fraction of irrelevant ones. Preparing such states is difficult, and it remains unclear how such a state preparation could take place without solving the electronic Schrödinger equation to high accuracy in the first place. Independently, two groups recently proposed methods with state-preparation protocols based on time-evolved or Hamiltonian-simulation QSCI, which could perhaps provide a workaround for this problem43, 44.
A second issue arises from the noisy nature of currently available quantum devices. Even though the target ansatz is, in principle, particle-conserving (i.e., has a desired number of $\alpha$ and $\beta$ electrons), the presence of hardware noise leads to many samples having the wrong number of particles. For example, in Ref. 33, only 0.44% of samples for the iron-sulfur cluster [2Fe-2S] (30e,20o) are within the target particle sector. Solutions to this problem have been proposed in the form of methods like the self-consistent configuration recovery (S-CORE)33, which seek to recover nearby bit-strings with a correct particle number from the set of erroneous bitstrings.
In our work, we will assume that these problems could be solved by a combination of improvements in quantum hardware and algorithms. Thus, we will not consider the state preparation or the noise problem but rather assume that an exact ansatz can be found. To this end, we rely on determinants sampled from CASCI wave functions. This approach eliminates complications related to noisy quantum devices, allowing us to focus on assessing the intrinsic efficiency and the quality of QSCI expansions. Ultimately, we demonstrate that even with an ideal sampling wave function in a noise-free environment, the stochastic nature of sampling itself either introduces significant inefficiencies or leads to less compact wave functions. Less compact CI expansions (i.e., more determinants for the same energy) lead to higher (classical) computational costs associated with the diagonalization of the CI Hamiltonian for the same final accuracy in the computation. These limitations undermine any practical advantage of QSCI over well-established classical SCI heuristics. Furthermore, enhancements like ext-SQD40, which add determinants post-sampling, could equally be applied to classical SCI methods, negating any unique benefit of quantum sampling.
2 Theory
A configuration interaction (CI) wave function can be expressed as a linear combination of Slater determinants, $\left|\Phi_{I}\right\rangle$,
$\left|\Psi\right\rangle=\sum_{I}c_{I}\left|\Phi_{I}\right\rangle,$ (1)
with the CI coefficients, $c_{I}$. The included determinants can be selected by excitation level out of a reference determinant (e.g., Hartree-Fock), forming the CISD, CISDT, etc. hierarchy45, which eventually terminates at the FCI expansion when $N$-tuple excitations are included. Alternatively, as is done in SCI expansions10, 11, 12, 13, a smaller set of important determinants can be selected based on classical heuristics. This approach recognizes that only a small fraction of determinants in the FCI expansion contributes significantly to the electronic wave function for most molecular systems 46.
The electronic Schrödinger equation is solved in the subspace spanned by the chosen Slater determinants via diagonalization of the CI Hamiltonian matrix
$H_{IJ}=\left\langle\Phi_{I}\left|\hat{H}\right|\Phi_{J}\right\rangle,$ (2)
where $H_{IJ}$ are matrix elements of the electronic Hamiltonian $\hat{H}$. The eigenvalues of the CI Hamiltonian matrix correspond to the electronic energy levels, and the eigenvectors (or CI vectors) contain the CI coefficients for the associated wave functions. The CI coefficients have a direct connection to the probability amplitudes for some particular determinant $\left|\Phi_{I}\right\rangle$, with the probability given by the norm-square of the CI coefficient, $|c_{I}|^{2}$. This probabilistic interpretation forms the basis of quantum sampling techniques, such as those used in QSCI, where determinants are stochastically selected by drawing samples from a wave function prepared on the quantum device.
In addition to QSCI, we will also consider a representative of classical SCI. In particular, we will consider Heat-bath Configuration Interaction12 (HCI). In HCI, important determinants for the CI expansion are added based on their expected contribution to the wave function. It uses a parameter $\varepsilon$ to screen a potential determinant $\left|\Phi_{I}\right\rangle$ and includes it if any of the existing determinants in the SCI expansion satisfy
$\left|H_{IJ}c_{J}\right|>\varepsilon.$ (3)
Here, $H_{IJ}$ is the Hamiltonian matrix element between an already included determinant $\left|\Phi_{J}\right\rangle$ and the candidate determinant $\left|\Phi_{I}\right\rangle$, while $c_{J}$ is the coefficient of the included determinant $\left|\Phi_{J}\right\rangle$. This ensures efficient selection by focusing on determinants with significant coupling, typically yielding compact and accurate CI expansions.
The HCI method works in an iterative fashion. Starting from some initial CI expansion (e.g., the HF determinant), the following steps are repeated: 1) the CI Hamiltonian is diagonalized, and the CI coefficients of the current determinant set are obtained; 2) determinants are added according to the criterion in eq. (3), with $\varepsilon$ as a parameter. These steps are repeated until no more determinants can be added according to eq. (3) (in the original work12, when the number of new determinants is less than 1% of the number of determinants already selected).
3 Computational Details
Hartree-Fock and CASCI (10e, 22o) wave functions, as well as one- and two-electron molecular orbital (MO) integrals (cc-pVDZ47 basis) of the nitrogen molecule, were obtained using PySCF 48. We considered structures near the equilibrium geometry ($R=1.09$Å) and along the potential energy surface $R=0.7$Å to $R=3.0$Å. Additionally, we investigated the iron-sulfur cluster [2Fe-2S] (30e,20o), where the one- and two-electron integrals were taken from the FCIDUMP file published by Li and Chan49, 50. Throughout the paper, we adopted the notation ($N$e, $M$o) for the active spaces, where $N$ is the number of active electrons and $M$ is the number of active orbitals.
The QSCI calculations were performed by drawing determinant samples from the CI vectors from the CASCI calculations with the probability of selecting a particular determinant given as $|c_{I}|^{2}$. The number of samples corresponds to the number of times the ansatz is measured. Our approach constitutes an ideal limit of QSCI since it assumes perfect state preparation of the exact CI vector and avoids any quantum device noise (i.e., non-ideal execution of the quantum circuits). The PyCI library51 was used to diagonalize the resulting CI Hamiltonian. HCI calculations were also performed using PyCI.
To assess the impact of an approximate ansatz in the sampling routine of QSCI, we obtained state vectors of equilibrium N2 for the unitary cluster Jastrow (UCJ)34 and local UCJ (LUCJ)35 ansatz using ffsim52. For both methods, two layers were used, and the ansatz parameters were assigned using the $t_{2}$ amplitudes of a classic CCSD simulation. This mimics the state preparation procedure used in recent works of IBM Quantum employing SQD/QSCI on quantum hardware33.
4 Results and Discussion
We evaluate the performance of the QSCI method on two distinct systems: the N2 molecule and the [2Fe-2S] iron-sulfur cluster. The N2 molecule is a prototypical example of a system dominated by weak, dynamic correlation, with static correlation effects eventually becoming significant as the triple bond is stretched. In contrast, the [2Fe-2S] cluster represents a more strongly correlated system, providing a challenging test for the method in capturing multi-reference effects. By comparing QSCI to classical heuristics like HCI for these cases, we can assess the efficiency and accuracy of the CI expansions produced by the QSCI method.
4.1 Potential energy curves of the nitrogen molecule
We first consider the ground state potential energy curve of N2/cc-pVDZ (10e, 22o). This active space is sufficiently small to allow CASCI reference calculations, which involve $0.693\times 10^{9}$ determinants in the CASCI expansion. We use the CI vector from the CASCI wave function to provide the QSCI samples. With this definition, QSCI will recover the full correlation energy from the reference CASCI wave function in the limit of infinite sampling.
As shown in Figure 1, we find a systematic convergence towards the reference CASCI energy when tightening the threshold $\varepsilon$ for the HCI wave functions. At $\varepsilon=10^{-3}$, the potential energy curve is in qualitative agreement with the reference CASCI result, and energy errors on the order of one milli-Hartree are reached with $\varepsilon=10^{-4}$. Similarly, the QSCI wavefunction converges systematically towards the reference CASCI result with an increasing number of samples. The QSCI wave function provides energy errors on the order of $10^{-2}$ Hartree with $10^{6}$ samples, which is quite similar to the HCI ($\varepsilon=10^{-3}$) result. Errors on the order of one milli-Hartree can be reached after using about $10^{8}$$10^{9}$ samples. Due to the probabilistic nature of the QSCI sampling, different runs will generally select different sets of determinants. Thus, when a small number of samples is used, there is a noticeable variance in the energies from different runs. This highlights that, although the process of solving the SCI Hamiltonian with some particular set of determinants leads to an energy estimation with no inherent statistical variance (unlike, e.g., VQE), the selection of the determinants themselves in the QSCI process is probabilistic, meaning that there is still statistical noise in the QSCI process. Fortunately, this variance becomes less noticeable when the number of samples is increased.
The rightmost panel of Figure 1 shows the number of determinants included in the CI expansion across the nitrogen potential energy surface. For both HCI and QSCI, the size of the CI expansion increases rapidly when tightening the threshold $\varepsilon$ (for HCI) or increasing the number of samples $N$ (for QSCI). Notably, we find that the HCI ($\varepsilon=10^{-3}$) result and QSCI with $10^{6}$ samples align quite well in terms of the number of determinants (and as mentioned before, in the energy). This suggests that, at least for this system, the quality of the determinants selected by QSCI is comparable to that obtained from classical heuristics (i.e., HCI), which is certainly encouraging. However, a sampling problem is also apparent. With $10^{6}$ samples, we find around $15\times 10^{3}$ determinants with QSCI. Increasing the number of samples to $10^{9}$, the number of determinants increases only to about $565\times 10^{3}$ determinants, which means that a 1000-fold increase in the sampling yields only a 38-fold increase in the number of found determinants. This happens because the same determinants are repeatedly sampled, providing no new information for the QSCI process and leading to a high number of “wasted” quantum samples, translating directly to immense measurement overheads. Coupled with the fact that obtaining good accuracy in the energy requires a significant number of determinants in the CI expansion, converging a QSCI wave function to high accuracy will be challenging.
Figure 1: Ground-state potential energy curve of N2/cc-pVDZ . The left panel shows total energies from HCI (solid lines), QSCI (dots), and CASCI (solid black line) wave functions with different thresholds $\varepsilon$ (HCI) and number of samples $N$ (QSCI). The middle panel shows energy errors relative to the CASCI reference. The right panel shows the number of determinants in the CI expansions. For the QSCI results, we include ten independent runs and indicate the variance across runs with error bars at a 95% confidence interval.
Next, we examine the sampling problem in more detail, focusing on N2 at $R=1.09$. In Figure 2, we show the energy and error relative to the CASCI energy for wave functions produced by QSCI and HCI, plotting the resulting energy against the number of determinants of the respective CI expansions. With HCI, each data point in the plot corresponds to a unique converged HCI calculation with increasingly tight $\varepsilon$, logarithmically spaced ($\varepsilon=10^{-1.0},10^{-1.1},10^{-1.2},...,10^{-6.0}$). Similarly, with QSCI, each point corresponds to a unique QSCI instance with an increasing number of samples, also logarithmically spaced ($10^{1.0},10^{1.1},...,10^{10.5}$). Clearly, the quality of HCI (blue lines) and QSCI (red dots) CI expansions are almost identical across a wide number of determinants, perhaps with a slim advantage to HCI.
Figure 2: Ground-state energy of the N2/cc-pVDZ molecule ($R=1.09$) with a (10e,22o) active space. The left panel shows total energies from HCI, QSCI, and CASCI wave functions as a function of the number of determinants included in the CI expansion. The middle panel shows energy errors relative to the CASCI reference. The right panel shows the determinants per sample for QSCI as a function of the number of samples.
However, as shown in the rightmost panel of Figure 2, the sampling issue of QSCI eventually becomes severe, with very few new determinants being discovered. We quantify this in terms of the number of determinants per sample (DPS). Initially (at $10^{3}$ samples), sampling from the CI vector yields about 0.1 determinants per sample. As more determinants are found, already-found determinants are frequently sampled again. As a result, the DPS drops to 0.01 at $10^{6}$ samples and 0.0005 at $10^{9}$ samples. As shown in Figure S2, if one wanted to target micro-Hartree precision on N2, it would require an estimated $4\times 10^{14}$ samples.
To understand this problem further, we consider results obtained by sampling from the scaled probability distribution
$\tilde{p}_{I}=\frac{(p_{I})^{\alpha}}{\left(\sum_{J}(p_{J})^{\alpha}\right)}$ (4)
where $\alpha$ is a scaling parameter and $p_{I}=|c_{I}|^{2}$ is the original CI vector probability of a given determinant $\left|\Phi_{I}\right\rangle$. Such a probability distribution allows us to tune the sampling by changing the weights of the determinants in the original CI vector, making it more likely (or less likely) to find new determinants depending on the value of $\alpha$. The original distribution of the determinants is recovered with $\alpha=1$. Letting $\alpha>1$ emphasizes the larger weights in the probability distribution, while letting $\alpha<1$ emphasizes smaller weights in the probability distribution. At $\alpha=0$, this distribution corresponds to uniform sampling from the correct particle sector (see also Figure S1).
Figure 3: Ground-state energy of the N2/cc-pVDZ molecule ($R=1.09$) with a (10e,22o) active space. The left panel shows energy errors from HCI and QSCI wave functions (relative to the CASCI reference) as a function of the number of determinants included in the CI expansion. The right panel shows the determinants per sample for QSCI as a function of the number of samples. For QSCI, we include sampling from the original CASCI distribution, as well as scaled CASCI distributions (see Eq. (4)), and from UCJ and LUCJ wave functions.
Such distributions are probably not realizable on real quantum hardware (except $\alpha=0$) but allow us to reason about the properties of alternative probability distributions while keeping the ranking of the CI weights intact (i.e., if $p_{I}>p_{J}$, the scaled distributions also satisfy $\tilde{p}_{I}>\tilde{p}_{J}$). In Figure 3, we show the effect of the scaled distribution for our N2 example. Letting $10\leq\alpha<$ 1 will increase the frequency of finding determinants with small weights, alleviating the sampling problem but making the resulting CI expansions less compact. Clearly, for $\alpha=0.5$ (light blue dots) in Figure 3, the sampling problem is essentially not present, with the DPS remaining high, even when going towards a large number of samples. The resulting CI expansion is, for a low number of determinants, not very good (note the high-error cyan points in the left panel of Figure 3). As the number of samples increases, most of the important determinants become selected, and a reasonably good energy is obtained, albeit with much less compact CI expansion, with around twice the number of determinants for the same energy accuracy as HCI or QSCI (with an unmodified probability distribution, $\alpha=1$). It should be noted that the (reduced) compactness of the wave function translates directly to the costs of the classical diagonalization step, with too-large expansions eventually becoming computationally intractable. In fact, we need so many determinants for $\alpha=0.5$ that the memory and compute requirements in the solution of the SCI Hamiltonian begin to become significant.
On the other hand, when the frequency of finding determinants with large weights is enhanced ($\alpha>1$), finding new determinants is difficult, but any selected determinants will most likely contribute significantly to the CI expansion. Thus, for $\alpha=2.0$ (see Figure 3, orange dots), we observe more severe sampling issues manifesting as very low DPS already at a low number of samples. However, the resulting CI expansions are very compact, even outperforming HCI by a slim margin. For example, using QSCI ($\alpha=2.0$), we obtain an energy error of 0.037 Hartree with 1050 determinants (from $2.5\times 10^{9}$ samples), where similar accuracy with HCI required 1100 determinants. Standard QSCI ($\alpha=1.0$) requires more determinants (around 1500), i.e., it is less compact but finds these determinants with only about 40,000 samples for a similar error. These results using artificially weighted distributions highlight an essential and fundamental issue with selecting determinants by sampling: the probability distribution can either be good at discovering many new determinants or good at selecting a compact CI expansion, but not both at the same time.
An additional factor to consider is that preparing an FCI-quality wave function on the quantum computer is difficult in the first place. In practice, the determinants would likely be sampled from a lower-quality wave function. We include examples of such wave functions in Figure 3, exemplified by the 2-UCJ34 and heavy-hex 2-LUCJ35 wave functions, with initial parameters assigned from CCSD $t_{2}$-amplitudes. This choice of ansatz has proven popular for previous hardware demonstrations33, although we point out that not a full 2-layer LUCJ was used in these works but rather a more approximated approach, resembling an ansatz structure between one and two layers. We find that sampling from the UCJ and LUCJ wave functions gives significantly worse results than the previously discussed sampling from the CASCI wave function. This manifests both in terms of a lower number of determinants per sample and significantly worse energies, even for relatively large CI expansions. The worse energies are likely due to a different ranking of the determinant weights, as shown in Figure S3. Curiously, LUCJ has a slightly better determinant selection than UCJ, although this is coupled with a more severe difficulty in finding new determinants and has an overall worse energy than QSCI sampled from the ideal CI vector. The less-than-stellar performance of the approximate wave functions (UCJ/LUCJ) could, at least in principle, be improved by actually optimizing the ansatz rather than just assigning parameters based on CCSD $t_{2}$ amplitudes. However, to the best of our knowledge, implementing this in practice proves to be a significant challenge and would still suffer the fundamental problems of QSCI reported above.
4.2 More correlated systems: Iron-sulfur clusters
Next, we consider the more challenging electronic structure of the iron-sulfur cluster [2Fe-2S] in a (30e,20o) active space. The one- and two-electron integrals for this system were obtained from Li and Chan49, 50. We note that the same system has been studied in several QSCI works33, 40. The number of determinants in this space ($240\times 10^{6}$) is small enough that a CASCI wave function can comfortably be obtained, so we will use this as the sampling wave function for the QSCI.
Figure 4: Convergence in the total energy for the HCI and QSCI approaches for the [2Fe-2S] system. The left panel shows the total energy as a function of selected determinants. The CASCI energy is given as a reference (dashed line). The CASCI wave function contains $240\times 10^{6}$ determinants. The middle panel plots the error in the energy relative to the CASCI reference. The right panel displays the number of determinants per sample for QSCI. For HCI, we include results starting from an HF-like determinant (blue lines, “bad guess”) or determinants with anti-ferromagnetically coupled spins on the iron centers (green lines, “good guess”).
Figure 4 shows the convergence in the energy as a function of the number of determinants included in the CI expansions for HCI and QSCI wave functions. For the HCI results, we include results for two different initial guesses. First, starting from an HF-like determinant (blue lines), which constitutes the simple yet “bad guess”. As the orbitals from Ref. 49 are local orbitals, the HF-like determinant corresponds to filling all electrons into the left side of the iron-sulfur cluster, which is clearly an unphysical configuration. In our second approach, we start from a set of $M_{s}=0$ determinants (green lines) with five spin-up electrons on one of the iron centers (Fe 3d orbitals) and five spin-down electrons on the other iron center. This corresponds to a configuration of anti-ferromagnetically coupled spins and would be a “good guess” based on a chemical understanding of the system. It also happens to be the largest weight in the CASCI expansion. The initial guess also impacts the magnitude of the “correlation energy”, defined as the energy difference between the single-determinant Hartree-Fock and the CASCI result. For the “bad guess”, one obtains a correlation energy of $-9.5$ Hartree, while for the chemically intuitive anti-ferromagnetically coupled set of determinants, one finds a smaller correlation energy of just $-0.65$ Hartree. In Figure 4, we show the HCI results and see that the HCI expansions are also sensitive to the choice of starting determinants, with chemical intuition allowing for more compact CI expansions. For those of us who lack good chemical intuition, the good news is that after including a sufficiently large number of determinants (after around $10^{5}$), the two HCI calculations attain comparable energies, even when a bad initial guess is used. With the bad HCI starting guess (blue lines), the QSCI wave function (red lines) is initially (i.e., at large energy errors) more compact than the HCI one until a cross-over point at around 10,000 determinants (at an energy error of about 0.2 Ha). Beyond this point, for the same accuracy result, the QSCI wave function always requires more determinants than the HCI expansions. The difference is even more stark when comparing QSCI with the good starting guess of HCI expansion (green lines). Here, the QSCI wave function is consistently less compact. For example, HCI manages an energy error below 0.1 Hartree with 4,260 determinants, which is more than one order of magnitude smaller than the 50,551 determinants required by QSCI for the same energy.
On the other hand, the problem of repeated sampling of already-seen determinants is less severe for the [2Fe-2S] system. Initially, almost every quantum sample leads to a new determinant (two, if the $\alpha$ and $\beta$-strings of a sample are not identical). As sampling progresses, the number of new determinants per sample begins to level off. As a result, at $10^{8}$ samples, only around $10^{7}$ determinants are found. The problem of repeated sampling is less pronounced compared to the nitrogen molecule since the CASCI wave function amplitudes have a rather flat and slowly decaying distribution with no single large weight (compare blue and black lines in Figure S4). This makes finding unseen determinants relatively easier. However, finding many determinants is not the primary goal of a selection algorithm for SCI. Rather, it is finding good determinants, which allow for a compact yet accurate CI expansion. As evidenced by the results in Figure 4, QSCI is not optimal in this respect and will thus need significantly more computational resources for the classical diagonalization step.
A careful reader may have noticed that our CI expansions (both HCI and QSCI) are much smaller than those reported in previous QSCI works (see, for example, Ref. 33 and Figure S10 therein). In Ref. 33, on the exact same [2Fe-2S] system, around $100\times 10^{6}$ determinants are required to reach an energy of around $-116.60$ Hartree, more than two orders of magnitude greater than the around $0.81\times 10^{6}$ determinants required in our calculations. The reportedly large number of determinants was especially surprising to us, considering that the entire CASCI space contains only $240\times 10^{6}$ determinants. This means that in some of the larger calculations of Ref. 33, which include nearly $200\times 10^{6}$ determinants, more than 80% of all determinants have been selected for the SCI expansion. We believe this discrepancy could be related to a slightly suboptimal SCI solver. A cursory glance at the implementation53 reveals that the approach relies on the kernel_fixed_space selected CI subroutine from PySCF48. As mentioned in the comments in the source code of the PySCF implementation, this subroutine is “…an inefficient dialect of Selected CI using the same structure as determinant based FCI algorithm”. In particular, this implementation does not select individual determinants. Instead, a set of $\alpha$- and $\beta$-strings are selected, and every determinant formed by all combinations thereof is included in the CI expansion. Such an algorithm may include up to quadratically more determinants in the CI expansion compared to individually selecting determinants. While this partially alleviates the sampling problem (since many more determinants are included), it somewhat contradicts the core aim of selected CI expansions – namely, to obtain good wave functions with as few determinants as possible.
5 Conclusion
In this work, we have explored the application of the Quantum Selected Configuration Interaction (QSCI) method in an ideal setting without device noise and assuming an optimized exact wave function ansatz is available. We investigated the performance and limitations of QSCI when applied to the nitrogen molecule (N2) and the iron-sulfur cluster [2Fe-2S] and identified two fundamental challenges that significantly impede the practicality of QSCI in realistic electronic structure calculations.
In the first case, exemplified by the nitrogen molecule, the QSCI method produces high-quality CI expansions comparable to those obtained through classical SCI heuristics. However, achieving high accuracy becomes increasingly difficult due to the problem of repeatedly sampling the same determinants. This inefficiency leads to difficulties in finding new determinants, making the process less effective as more samples are added. Since achieving near-FCI quality results (i.e., recovering the correlation energy) requires CI expansions with a large number of determinants, the high sampling demands of the QSCI method make it impractical for useful applications.
Alternatively, one may find systems such as the more strongly correlated iron-sulfur cluster [2Fe-2S], where sampling turns out to be less of a problem due to a more flat distribution of the CI coefficients. Unfortunately, the CI expansions generated by the QSCI method are much less compact than those generated by classical heuristics in these cases. HCI attains better energies with fewer determinants, leading to lower costs in the expensive diagonalization step compared to QSCI. This, again, limits the practical usefulness of QSCI.
Our investigation into scaled probability distributions further supports the idea that the sampling problem is fundamental to QSCI. We demonstrate that tuning the probability distribution can improve either the discovery of new determinants or the compactness of the CI expansion, but not both simultaneously. This trade-off highlights a key limitation of the QSCI approach. We also find that these problems significantly increase when using an approximated ansatz, such as LUCJ, for the sampling routine.
Overall, while QSCI can, in principle, be used to generate selected CI expansions, its practical utility is hindered by its reliance on large numbers of samples or its difficulty in producing compact wavefunctions. Hence, even with the perfect quantum settings assumed for QSCI in this work, it ultimately falls behind more effective classical SCI heuristics.
{acknowledgement}
We acknowledge the financial support of the Novo Nordisk Foundation for the focused research project Hybrid Quantum Chemistry on Hybrid Quantum Computers (HQC)2, grant number NNFSA220080996. Computations and simulations for the work described herein were supported by the DeiC National HPC Centre, SDU.
{suppinfo}
Effect of probability density function scaling (Eq. (4)), N2 ground-state energy error as a function of the number of quantum samples, and CI probability distribution plots for N2 and [2Fe-2S] systems.